OpenStudy (anonymous):

-5/7+7/12 write in simplest form

OpenStudy (ybarrap):

-5/7 + 7/12 = (-5*12 + 7 *7 )/ (7*12) = (-60 + 49)/86 = -11/86 Ans -11/86